PHYSICS HMWK:

Determine the energy required to accelerate an electron between each of the following speeds.

(a) 0.500c to 0.895c
MeV

(b) 0.895c to 0.940c
MeV
Answer #1

a)

[(1/(1-v1^2/c^2))^.5 - (1/(1-v2^2/c^2))^.5] x Mo x c^2
= [(1/(1-(0.895c)^2/ c^2))^.5) - [(1/(1-(0.500c)^2/ c^2))^.5) x Mo x c^2
= (1/(1-.801))^.5 - (1/(1-.250))^.5] x Mo x c^2
= [2.148 - 1.55] x Mo x c^2
= .993 x Mo x c^2 = .993 x 9.1096 x 10^-31 kg x (3 x 10^8 m/s)^2
= 8.14 x 10^-14 kg m^2/s^2 = 8.15 x 10^-14 J

b)

same equation...
= [(1/(1-(0.940c)^2/ c^2))^.5) - [(1/(1-(0.895c)^2/ c^2))^.5) x Mo x c^2 = = 6.42 x 10 ^-14 J